OpenStudy (anonymous):

Hw to integrate x^2.e^x^(3-1)

OpenStudy (anonymous):

Do i use product rule

hartnn (hartnn):

is the question \(\huge \int x^2 e^{x^3-1}dx\) ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

then do a simple 'u' substitution,no need of integration by parts

hartnn (hartnn):

\(\Large u = x^3-1\) du =... ?

OpenStudy (anonymous):

3x

hartnn (hartnn):

just 3x ? u = x^3-1 differentiate w.r.t x, du/dx =... ?

OpenStudy (anonymous):

3x^2

hartnn (hartnn):

correct so, du/dx = 3x^2 but in the question we have x^2 dx

OpenStudy (anonymous):

Yea

hartnn (hartnn):

so, du = 3x^2 dx x^2 dx = du/3 got this ?

OpenStudy (anonymous):

Oh yes

hartnn (hartnn):

now whats your new integral, only in 'u' ?

OpenStudy (anonymous):

Why differentiating when asked to integerate

hartnn (hartnn):

we are changing the variable from 'x' to 'u' so we need to find du in terms of dx to find du, we need to differentiate u

hartnn (hartnn):

while doing substitution, we need to completely change the function, including dx

OpenStudy (anonymous):

X^2. e^(3x^2-1)

hartnn (hartnn):

? there should be no 'x' in your new function

hartnn (hartnn):